If the length of a garden is 6 meters longer than the width and the area is 40m squared, how many meters of fencing need to be purchased?
Answer by ankor@dixie-net.com(22740)

If the length of a garden is 6 meters longer than the width and the area is 40m squared, how many meters of fencing need to be purchased?
:
Find the width from the area: L * W = 40 sq/m
W(W+6) = 40
W^2 + 6W - 40 = 0
Factors to:
(W+10)(W-4) = 0
The positive solution is what we want here
W = 4 m is the width
then
L = 4 + 6
L = 10 m is length
:
Find the amt of fencing needed by finding perimeter
2(10) + 2(4) = 28 m of fencing required